1. Calcium metal can be obtained by the direct electrolysis of molten CaCl2, at a voltage of 3.2 V.
  2. How many joules of electrical energy are required to obtain 12.0 lb. of calcium?
  3. What is the cost of the electrical energy obtained in (a) if electrical energy is sold at the rate of nine cents per kilowatt hour?
1
Expert's answer
2021-02-11T04:29:24-0500

2. 12 lb = 5443,11 g.

m=ItE/96500

I=mx96500/Et= 5443,11 x 1608/20x1=

364 688,37.

E=I x V= 364 688,37 x 3.2/1000 = 1 167,00278 kwat

3. Cost for 1 kwat x 1 167,00278.